# g-DPO: Scalable Preference Optimization for Protein Language Models

### v1

@@ Abstract
Direct Preference Optimization (DPO) is an effective approach for aligning protein language models with experimental design goals. However, DPO faces a scalability bottleneck: the number of possible training pairs grows quadratically with the number of labeled sequences, leading to prohibitive training times even for modestly sized datasets. We introduce g-DPO, a framework that (i) uses sequence space clustering to prune redundant pairs while preserving training signal, and (ii) amortizes likelihood computations with group-based approximations. Across three protein engineering tasks, g-DPO maintains in-silico and in-vitro performance that is statistically indistinguishable from standard DPO, while converging $1.8$ to $3.7$ times faster, with greater gains expected as the size of the dataset increases.
@@ 1 Introduction
We introduce group-DPO (g-DPO), a DPO framework for experimentally labeled protein data that addresses these challenges. This study contributes: (1) Scalable preference sampling: Sequence-space clustering prunes redundant comparisons and focuses training on informative local neighborhoods in input space. (2) Efficient training: Building on this clustering, we exploit union masking to amortize log-likelihood computations across groups of sequences, further improving convergence time. (3) Empirical validation: Across three protein mutational landscapes, g-DPO maintains the in-silico and in-vitro performance of standard DPO, while converging $1.8$ to $3.7$ times faster, with larger convergence time gains expected as the size of the dataset increases.
@@ 3 Methods: Group-DPO (g-DPO)
In NLP, preferences are inherently structured: multiple outputs are generated from the same prompt, and annotators decide which is better [22, 3]. This provides a shared context that makes likelihood ratios between outputs meaningful. On the other hand, comparing outputs from different prompts yields likelihoods ratios that are dominated by prompt differences rather than output preference. Proteins have a related dynamic: even when variants come from the same wild type, comparing distant variants conflates the effects of many mutations, collapsing into coarse signals, whereas local comparisons better capture subtle, non-additive effects [12, 10]. Our work, therefore, leverages proximity in sequence space to construct preference pairs, enabling the model to learn fine-grained mutational effects while reducing the number of training pairs.
@@ 3.2 Group sampling
[…]
@@ 4 Experiments
Figure 2: In-silico evaluation of g-DPO. (a) Spearman correlation ($\rho$) of evo-tuned, DPO, and g-DPO models across the three datasets. (b) Relative convergence speed of g-DPO compared to DPO, measured on a single NVIDIA A100 GPU, with $\Delta\rho$ indicating performance differences. (c) Predicted property distributions of sequences generated from the reference, DPO, and g-DPO models using beam search. Kolmogorov-Smirnov tests confirm that g-DPO and DPO yield nearly identical, and improved distributions over the reference.
Evaluating generative models for protein optimization is challenging and ultimately requires experimental evidence. Our goal is to test whether g-DPO (i) preserves or improves model quality relative to DPO while (ii) reducing training cost, and (iii) whether in-silico improvements translate to experimental outcomes. We evaluate g-DPO on experimentally measured variants derived from three wild type proteins:
  • Anti-SARS-CoV-2 VHH [9]: thermostability ($N=462$, mutation breadth 47.1%).
  • Trastuzumab scFv: expression ($N=76$, mutation breadth 13.1%).
  • Haloalkane dehalogenase (DhaA): thermostability ($N=474$, mutation breadth 40.3%).
We use ESM-2-650 [14] as our pre-trained model. Additional dataset and training details are provided in Appendix C.
4.1 In-silico validation
For each dataset, we report (i) rank correlation between PLL scores and experimental measurements on holdout test sets, which tests whether model likelihoods align with ground truth preferences; (ii) convergence time to early stopping, to measure training efficiency; and (iii) generative quality, measured by scoring sequences generated with beam search [31]. We score the generated sequences with independent predictors trained on the same experimental datasets.
Both DPO and g-DPO improve rank correlation relative to the evo-tuned reference model (Figure 2a) and shift the distribution of generated sequences toward better predicted function (Figure 2c).
@@ 4.1 In-silico validation
Figure 3: In-vitro validation. Distribution of assay measurements for sequences designed with g-DPO and DPO on (a) thermostability of DhaA and (b) expression of Trastuzumab. Distributions show that both models yield comparable outcomes.
The distributions produced by DPO and g-DPO are statistically indistinguishable, indicating that g-DPO preserves generative quality while reducing training cost (Figure 2b). Statistical significance was assessed with two-sample KS tests, with full results reported in Table 2.
4.2 In-vitro validation
To test whether in-silico improvements translate to experimental outcomes, we validated a subset of generated sequences in the lab for two of the three optimization tasks: thermostability of DhaA and expression of Trastuzumab. Candidates were down-selected from the in-silico generation pool using a criterion that averages performance across the top-$k$ predicted sequences ($k=3$), following the Monte Carlo sampling strategy of DiscoBAX [16]. Assay results confirm that DPO and g-DPO yield comparable experimental outcomes, with no significant differences observed between their distributions.
@@ 5 Conclusion, limitations and future work
We present g-DPO, a scalable variant of DPO for PLMs, that combines sequence-space clustering, to prune redundant preference pairs, with grouped likelihood amortization, to reduce computation. Across three protein engineering tasks, g-DPO matches DPO performance on both in-silico metrics and in-vitro results, while converging up to 3.7× faster. In practice, moderate clustering with grouping yields the best trade-off; excessive clustering removes training signal, and grouping alone can degrade performance when mutation span is large and likelihood approximations break down.
@@ C.1 Training details
We start with pre-trained ESM-2-650 (weights are open-sourced) [14], which we fine-tune on evolutionarily related sequences of our wild type sequence for each corresponding dataset [1, 8]. We follow the evo-tuning framework defined by [1]. Then, this model is further fine-tuned using the g-DPO framework with SGD using a learning rate of $7\times 10^{-4}$ and no weight decay, with a 300-step linear warmup. The DPO loss is computed with $\beta=0.04$ with a batch size of 64. Sequences were clustered with a maximum union mask size of $0.3L$, and groups of $g=4$ were sampled from each cluster. The loss was applied over all within-group pairs. This was the final configuration, though we ran hyperparameter sweeps and expect the optimal group size to depend on dataset scale and clustering. Validation loss was monitored every 250 steps, with checkpointing on validation loss and relative early stopping on validation loss that checks there has not been more than a 1% improvement with a patience of 3 validation intervals. Training was performed on a single NVIDIA A100 GPU.

### v2

@@ Abstract
Direct Preference Optimization (DPO) is an effective approach for aligning protein language models with experimental design goals. However, DPO faces a scalability bottleneck: the number of possible training pairs grows quadratically with the number of labeled sequences, leading to prohibitive training times even for modestly sized datasets. We introduce g-DPO, a framework that (i) uses sequence space clustering to prune redundant pairs while preserving training signal, and (ii) amortizes likelihood computations with group-based approximations. Across three protein engineering tasks, g-DPO maintains in silico and in vitro performance that is statistically indistinguishable from standard DPO, while converging $1.7\times$ to $5.4\times$ times faster, with speedups that scale with dataset size and the structure of the underlying mutational landscape.
@@ 1 Introduction
We introduce group-DPO (g-DPO), a DPO framework for experimentally labeled protein data that addresses these challenges. This study contributes: (1) Scalable preference sampling: Sequence-space clustering prunes redundant comparisons and focuses training on informative local neighborhoods in input space. (2) Efficient training: Building on this clustering, we exploit union masking to amortize log-likelihood computations across groups of sequences, further improving convergence time. (3) Empirical validation: Across three protein mutational landscapes, g-DPO maintains the in silico and in vitro performance of standard DPO, while converging $1.7$ to $5.4$ times faster, with larger convergence time gains expected as the size of the dataset increases.
@@ 3 Methods: Group-DPO (g-DPO)
In NLP, preferences are inherently structured: multiple outputs are generated from the same prompt, and annotators decide which is better [22, 3]. This provides a shared context that makes likelihood ratios between outputs meaningful. On the other hand, comparing outputs from different prompts yields likelihoods ratios that are dominated by prompt differences rather than output preference. Proteins have a related dynamic: even when variants come from the same wild type, comparing distant variants conflates the effects of many mutations, collapsing into coarse preference signals, whereas local comparisons better capture subtle, non-additive effects [12, 10]. Our work, therefore, leverages proximity in sequence space to construct preference pairs, enabling the model to learn fine-grained mutational effects while reducing the number of training pairs.
@@ 3.2 Group sampling
Evaluating the DPO loss on a pair $(y_{w},y_{l})$ requires likelihoods for both sequences. For masked language models (MLMs) like ESM-2 [14], computing pseudo log-likelihood (PLL) [27] scores has become the standard way to approximate sequence likelihoods. This requires one forward pass per position and is therefore expensive. To reduce this cost, we exploit the union mask $D=M(\{y_{w},y_{l}\})$, and create a jointly masked input $y_{\setminus D}$, where all differing positions between $y_{w}$ and $y_{l}$ are masked. One forward pass on $y_{\setminus D}$ provides logits for all positions. We then approximate $\log p(y_{w})\approx\sum_{i\in D}\log p({y_{w}}_{i}|y_{\setminus D})$ (same for $y_{l})$. Positions outside $D$ do not need to be evaluated since their logits agree. This is equivalent to a mean-field approximation, where we assume that, conditional on the observed tokens, the masked positions are independent. Although approximate, it is reliable when the differing positions are few compared to the sequence length, since the error scales linearly with the size of the difference mask. Similar multi-mask approximations have already been applied successfully in prior work [19, 5, 21, 39].
@@ 4 Experiments
Figure 2: In silico evaluation of g-DPO. (a) Spearman correlation ($\rho$) of evo-tuned, DPO, and g-DPO models. (b) Convergence speedup of g-DPO vs. DPO (higher better), measured on a single NVIDIA A100 GPU; wall-clock times are in Table 2. (c) Predicted property distributions of sequences generated from the evo-tuned, DPO, and g-DPO models using beam search. KS tests confirm that g-DPO and DPO yield nearly identical, and improved distributions over the reference.
Evaluating generative models for protein optimization is challenging and ultimately requires in vitro validation. We test whether g-DPO (i) preserves or improves model quality relative to DPO, (ii) reduces training cost, and (iii) whether in silico gains translate to experimental outcomes. We evaluate g-DPO on experimentally measured variants from three wild-type proteins:
Dataset Function NN Positional coverage
Anti-SARS-CoV-2 VHH Thermostability 462 47.1%
Anti-SARS-CoV-2 VHH L Thermostability 1833 92.4%
Trastuzumab scFv Expression 76 13.1%
Haloalkane dehalogenase (DhaA) Thermostability 474 40.3%
For VHH L we report training metrics and one in silico metric, as it is included for scaling analysis. We use ESM-2-650M [14] as our pre-trained model; further details are provided in Appendix C.
4.1 In silico validation
For each dataset, we report (i) rank correlation between PLL scores and experimental measurements on holdout test sets, which tests whether model likelihoods align with ground truth preferences; (ii) convergence time to early stopping, to measure training efficiency; and (iii) generative quality, measured by scoring sequences generated with beam search [31].
@@ 4.1 In silico validation
Figure 3: In vitro validation. Distribution of assay measurements for sequences designed with g-DPO and DPO on (a) thermostability of DhaA and (b) expression of Trastuzumab. Distributions show that both models yield comparable outcomes.
We score the generated sequences with independent predictors trained on the same experimental datasets.
Both DPO and g-DPO improve rank correlation relative to the evo-tuned reference model (Figure 2a) and shift the distribution of generated sequences toward better predicted function (Figure 2c). The distributions produced by DPO and g-DPO are statistically indistinguishable, indicating that g-DPO preserves generative quality while reducing training cost (Figure 2b). Statistical significance was assessed with two-sample KS tests, with full results reported in Table 3.
4.2 In vitro validation
To test whether in silico improvements translate to experimental outcomes, we validated a subset of generated sequences in the lab for two optimization tasks: thermostability of DhaA and expression of Trastuzumab. Candidates were down-selected from the in silico pool using a criterion that averages performance across the top-$k$ predicted sequences ($k=3$), following the Monte Carlo strategy of DiscoBAX [16]. Assay results confirm that DPO and g-DPO yield comparable outcomes, with no significant differences between their distributions.
@@ 5 Conclusion, limitations and future work
We present g-DPO, a scalable variant of DPO for PLMs, that combines sequence-space clustering, to prune redundant preference pairs, with grouped likelihood amortization, to reduce computation. Across three protein engineering tasks, g-DPO matches DPO performance on both in silico metrics and in vitro results, while converging up to 5.4× faster. In practice, moderate clustering with grouping yields the best trade-off; excessive clustering removes training signal, and grouping alone can degrade performance when mutation span is large and likelihood approximations break down.
@@ C.1 Training details
We start with pre-trained ESM-2-650M (weights are open-sourced) [14], which we fine-tune on evolutionarily related sequences retrieved via MMseqs2 [29] searches against ColabFold databases [mirdita2022colabfold]. We follow the evo-tuning framework defined by [1]. Then, this model is further fine-tuned using the g-DPO framework with SGD using a learning rate of $7\times 10^{-4}$ and no weight decay, with a 300-step linear warmup. The DPO loss is computed with $\beta=0.04$ with a batch size of 64. Sequences were clustered with a maximum union mask size of $0.3L$, and groups of $g=4$ were sampled from each cluster. The loss was applied over all within-group pairs. This was the final configuration, though we ran hyperparameter sweeps and expect the optimal group size to depend on dataset scale and clustering. Validation loss was monitored every 250 steps, with checkpointing on validation loss and relative early stopping on validation loss that checks there has not been more than a 1% improvement with a patience of 3 validation intervals. Training was performed on a single NVIDIA A100 GPU.
@@ D.1 In silico results
Dataset Training Time Total Run Time Speedup ($\times$)
g-DPO DPO g-DPO DPO Training Total
[…]
Table 2: Runtime comparison between g-DPO and DPO across datasets. Total run time includes training, clustering, and constant overhead. Training time is wall clock times for convergence. Speedup is computed as DPO / g-DPO.